The benches of a gallery in a cricket stadium are 1 m wide and 1 m high. A batsman strikes the ball at a level one metre above the ground and hits a mammoth sixer. The ball starts at 35 m/s at an angle of 53° with the horizontal. The benches are perpendicular to the plane of motion and the first bench is 110 m from the batsman. On which bench will the ball hit?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. θ = 53°, so cos 53° = 3/5 Sec2 θ = 25/9 and tan θ = 4/3 Suppose the ball lands on nth bench So, y = (n – 1)1 …(1) [ball starting point 1 m above ground] Again y = x tan θ – (gx^2 Sec^2 θ)/(2u^2 ) [x = 110 + n – 1 = 110 + y] ⇒ y = (110 + y)(4/3) – (10 (110+y)^2 (25/9))/(2 x 〖35〗^2 ) ⇒ 440/3 + 4/3 y - (25 (110+y )^2)/(18 x 〖35〗^2 ) From the equation, y can be calculated. ∴ y = 5 ⇒ n – 1 = 5 ⇒ n = 6. The ball will drop in sixth bench.